# Quasi one dimensional anomalous (rogue) waves in multidimensional nonlinear Schrödinger equations 1: fission and fusion

### v1

@@ Abstract
In this paper we study the first nonlinear stage of modulation instability (NLSMI) of $x$-periodic AWs in multidimensional generalizations of the focusing nonlinear Schrödinger (NLS) equation, like the non-integrable elliptic and hyperbolic NLS equations in $2+1$ and $3+1$ dimensions. In the quasi one-dimensional (Q1D) regime, where the wavelength in the $x$ direction of propagation is significantly smaller than in the transversal directions, the behavior is universal, independent of the particular model at leading order, and described by adiabatic deformations of the Akhmediev breather solution of NLS. Varying the initial data, the first NLSMI shows various combinations of basic processes like AW growth from the unstable background, followed by fission in the slowly varying transversal directions, and the inverse process of fusion, followed by AW decay to the background. Fission and fusion are critical processes showing similarities with multidimensional wave breaking, and with phase transitions of second kind and critical exponent $1/2$. In $3+1$ dimensions with radial symmetry in the transversal slowly varying plane, fission consists in the formation of an opening smoke ring centered on the $x$ axis. In the long wave limit, the Q1D Akhmediev breather reduces to the Q1D analogue of the Peregrine instanton, rationally localized in space. Numerical experiments on the hyperbolic NLS equation show that the process of "AW growth + fission" is not restricted to the Q1D regime, extending to a finite area of the modulation instability domain. The universality of these processes suggests their observability in natural phenomena related to AWs in contexts such as water waves, nonlinear optics, and plasma physics.
@@ 1 Introduction
rationally localized in space-time over $u_{0}$, together with their multi-mode generalizations [35], [25]. Peregrine and Akhmediev AWs are the simplest exact solutions describing the basic features of AWs in $1+1$ dimensions: an exponential growth, due to MI, from the unstable background $u_{0}(t)$, reaching the amplitude maximum in $(x_{0},t_{0})$, and the subsequent decay to the background, and they have been observed in many experiments; see, for instance: [13, 38, 68, 39, 48, 55].
In the NLS Cauchy problem of AWs for localized initial perturbations of the background $u_{0}(t)$, slowly modulated periodic oscillations described by the elliptic solution of (1) play a relevant role in the longtime regime [9, 10]. If the perturbations are periodic with period $L$, the finite-gap method [50, 26, 34, 41, 47, 40] has been adapted in [27, 29] to solve at leading order and in terms of elementary functions the NLS Cauchy problem of the AWs, in the case of a finite number $N=\lfloor\frac{L}{\pi}\rfloor$ of unstable modes. In the simplest case of a single unstable mode $k=2\pi/L$ ($N=1,\ \pi<L<2\pi$), one obtains an ideal Fermi-Pasta-Ulam-Tsingou (FPUT) recurrence of ABs, giving a quantitative description in terms of elementary functions of the AW recurrence observed in real and numerical experiments [66, 68, 39, 48, 55]. This one-mode FPUT recurrence can also be described by a simpler matched asymptotic expansion (MAE) approach [28], not requiring the powerful but more technical tools of integrability. See also [30] for a finite-gap model describing the numerical instabilities of the AB, and [31] for the analytic study of the linear, nonlinear, and orbital instabilities of the AB within the NLS dynamics. See [32] for the analytic study of the phase resonances in the AW recurrence, and see [58], [18] and [20] for the analytic study of the FPUT AW recurrence in other NLS type models: respectively the PT-symmetric NLS equation [3], the Ablowitz-Ladik lattice [2], and the massive Thirring model [64]. To explain quantitatively the $O(1)$ effects of a small loss or gain on the NLS AW recurrence observed in real and numerical experiments [39, 61], a finite gap perturbation theory of AWs has been successfully constructed in [14] (see also [15]), and then applied in [17] to the Ginsburg-Landau [49] and Lugiato-Lefevre [42] models, viewed as perturbations of NLS, and in [19] to hamiltonian and non hamiltonian perturbations of the Ablowitz-Ladik lattice.
@@
or, equivalently, the dependence on the extra space variables $y,z,\dots$ is slow. In this regime all the above MNLS equations are close to NLS, and one expects that suitable adiabatic deformations of the AB (2) and of the Peregrine instanton (3) be relevant. The goal of this paper is to uncover the universal analytic features of AWs of MNLS equations in the Q1D regime, concentrating our attentions on the first nonlinear stage of MI (NLSMI), and showing, at leading order, a universal behavior, independent of the particular model, and described in terms of suitable adiabatic deformations of the AB (2). Varying the initial data, the first NLSMI consists in various combinations of the following two basic processes. P1: the AW growth from the unstable background, followed by fission in the transversal slowly varying directions, with infinite speed at fission time. P2: the inverse process of fusion in the transversal slowly varying directions, with infinite speed at fusion time, followed by AW decay to the background. Fission and fusion are critical processes showing similarities to multidimensional wave breaking, and to phase transitions of second kind and critical exponent $1/2$. They may or may not coexist in the same nonlinear stage of MI, but while the universal process of “AW growth + fission” always occurs, fusion is less generic. We also show that, in the natural long wave limit, the above Q1D Akhmediev breather reduces to the Q1D analogue of the Peregrine AW, rationally localized in the space variables, also undergoing the above critical processes. At last we show, through numerical experiments on the hyperbolic NLS equation, that the process of “AW growth + fission” is by no means restricted to the Q1D regime, extending to a finite region of the MI domain of the equation (see [21] for more details). The recurrence of Q1D AWs, another important feature of $x$-periodicity, turns out to be different for different MNLS models, and the quantitative description of these differences is postponed to the subsequent paper [23]. Due to the universality of the above processes, it is very plausible that they be observable in natural phenomena in which AWs of multidimensional NLS equations be relevant, like in water waves, nonlinear optics, and plasma physics.
@@ 2 AW fission and fusion
$\begin{array}[]{l}u_{lin}(x,Y,t)=e^{2it}\Big{[}1+{\epsilon}\Big{(}\frac{2|\alpha(Y)|}{\sigma(k_{x})}e^{\sigma(k_{x})t+i\phi}\cos[k_{x}(x-x_{1}(Y))]\\ +\frac{2|\beta(Y)|}{\sigma(k_{x})}e^{-\sigma(k_{x})t-i\phi}\cos[k_{x}(x-\tilde{x}_{1}(Y))]\Big{)}+O({\epsilon})\mbox{ oscillations}\Big{)}\Big{]},\end{array}$ (8)
@@
$\begin{array}[]{l}x_{1}(Y)=\frac{\arg(\alpha(Y))+\pi/2}{k_{x}},\ \ \tilde{x}_{1}(Y)=\frac{-\arg(\beta(Y))+\pi/2}{k_{x}},\\ \alpha(Y)=e^{-i\phi_{1}}c^{*}_{+}(Y)-e^{i\phi_{1}}c_{-}(Y),\ \ \beta(Y)=e^{i\phi_{1}}c^{*}_{-}(Y)-e^{-i\phi_{1}}c_{+}(Y).\end{array}$ (9)
@@
matching successfully with the following adiabatic deformation of the AB (2) [28] when $t=O(\log\left(\frac{1}{{\epsilon}}\right))$
$\begin{array}[]{l}u_{1}(x,Y,t)={\cal A}\left(x-x_{1}(Y),t-t_{1}(Y),\phi\right)e^{2it+2i\phi},\\ t_{1}(Y)=\frac{1}{\sigma(k_{x})}\log\left(\frac{\sigma^{2}}{2{\epsilon}|\alpha(Y)|}\right).\end{array}$ (11)
Equation (11) describes at leading order the first NLSMI, the same for equations (4) and for all MNLS equations, up to small corrections whose analytic investigation is postponed to a subsequent paper. Therefore it describes the universal features of Q1D AWs. Also for the ENLS equation, well known to lead to a blow up of generic initial perturbations of $u_{0}$ [63], the first NLSMI is described by (11) if the MI time scale $\ln(1/{\epsilon})$ is much smaller than the Q1D scale $1/\delta$, implying that blow up occurs only after a number of recurrences.
To understand the main properties of the first NLSMI (11), we restrict the family of initial data (7) to
@@
Since $\dot{x}_{1}(0)=\dot{t}_{1}(0)=0$, and $\ddot{t}_{1}(0)>0$, it follows that, at the max point $(x,Y)=(x_{10},0)$, ${u_{1}}_{Y}={u_{1}}_{x}=0$, and ${u_{1}}_{YY}(x_{1}(0),0,t)=-\ddot{t}_{1}(0){\cal A}_{t}(0,t-t_{0})$, $|t-t_{0}|\ll 1$. Consequently $(|u_{1}(x_{10},0,t)|^{2})_{YY}=-\ddot{t}_{1}(0)\rho_{t}(t)$, implying that $\mbox{sign}\!\left((|u_{1}(x_{10},0,t)|^{2})_{YY}\right)=\mbox{sign}(t-t_{0})$, $|t-t_{0}|\ll 1$. It follows that, when the Q1D AW reaches it max, its $Y$-curvature changes sign, from negative to positive sign, corresponding to fission in the $Y$ direction. Fission and fusion as critical processes. Fission and fusion are critical processes showing similarities to multidimensional wave breaking according to the integrable dispersionless Kadomtsev-Petviashvili (dKP) equation $(u_{t}+\\ uu_{x})_{x}+u_{yy}=0$ (describing for instance the evolution of small amplitude, nearly one-dimensional waves in shallow water near the shore), and to phase transitions of second kind and critical exponent $1/2$. Indeed, i) in the dKP wave breaking, a smooth localized initial datum undergoes the first wave breaking in a space point $(x_{b},y_{b})$ at time $t_{b}$, constructed from the initial data via the spectral transform developed for integrable dispersionless PDEs [44, 45]. For $0<t-t_{b}\ll 1$ the multivaluedness region develops from the breaking point $(x_{b},y_{b})$ with the law $\Delta x=O\left((t-t_{b})^{3/2}\right)$ and velocity $\Delta\dot{x}=O\left((t-t_{b})^{1/2}\right)$ (zero speed at breaking time) in the $x$ direction, like in the $1+1$ dimensional reduction $u_{t}+uu_{x}=0$ (the Hopf equation), while it develops with the law $\Delta y=O\left((t-t_{b})^{1/2}\right)$ and velocity $\Delta\dot{y}=O\left((t-t_{b})^{-1/2}\right)$ (infinite speed at breaking time) in the slowly varying direction (see Figure 4) [45, 46]. ii) in the Landau-Ginzburg theory of phase transitions, the behavior of the order parameter around the critical temperature $T_{c}$ is a two-valued function proportional to $\pm\theta(T_{c}-T)\sqrt{T_{c}-T}$, where $\theta(\cdot)$ is the step function, and the susceptivity is singular at $T_{c}$, with critical exponent $1/2$ [52].
@@
Figure 7: The central figure shows the MI domain $0<k^{2}_{x}-k^{2}_{y}<4$ of the HNLS equation. The right figure shows four snapshots of the HNLS numerical evolution of a doubly periodic initial condition $1+{\epsilon}(e^{i\phi}\cos(k_{x}x)\cos(k_{y}y))$, with periods $L_{x}=4,\ L_{y}=10$, where $k_{x}=2\pi/L_{x},\ k_{y}=2\pi/L_{y}$ are the coordinates of the red points, and $\phi=\arccos(\sqrt{k^{2}_{x}-k^{2}_{y}}/2)$; this AW undergoes fission. The left figure shows four snapshots of the HNLS numerical evolution for the same initial condition, but with different periods $L_{x}=2.2,\ L_{y}=2.89$, corresponding to the green points; this AW does not undergo fission.
@@
Acknowledgments. F.C and P.M.S. have been supported by the Research Project of National Interest PRIN2020 No. 2020X4T57A; F.C. was also supported by PRIN2022 No. 20223T577Z.

### v2

@@ Abstract
In this paper we study the first nonlinear stage of modulation instability (NLSMI) of $x$-periodic anomalous waves (AWs) in multidimensional generalizations of the focusing nonlinear Schrödinger (NLS) equation, like the non-integrable elliptic and hyperbolic NLS equations in $2+1$ and $3+1$ dimensions. In the quasi one-dimensional (Q1D) regime, where the wavelength in the $x$ direction of propagation is significantly smaller than in the transversal directions, the behavior at leading order is universal, independent of the particular model, and described by adiabatic deformations of the Akhmediev breather solution of NLS. Varying the initial data, the first NLSMI shows various combinations of basic processes, like AW growth from the unstable background, followed by fission in the slowly varying transversal directions, and the inverse process of fusion, followed by AW decay to the background. Fission and fusion are critical processes showing similarities with multidimensional wave breaking, and with phase transitions of second kind and critical exponent $1/2$. In $3+1$ dimensions with radial symmetry in the transversal slowly varying plane, fission consists in the formation of an opening smoke ring centered on the $x$ axis. In the long wave limit, the Q1D Akhmediev breather reduces to the Q1D analogue of the Peregrine instanton, rationally localized in space. Numerical experiments on the hyperbolic NLS equation show that the process of "AW growth + fission" is not restricted to the Q1D regime, extending to a finite area of the modulation instability domain. The universality of these processes suggests their observability in natural phenomena related to AWs in contexts such as water waves, nonlinear optics, plasma physics, and Bose-Einstein condensates.
@@ 1 Introduction
rationally localized in space-time over $u_{0}$, together with their multi-mode generalizations [37], [27]. Peregrine and Akhmediev AWs are the simplest exact solutions describing the basic features of AWs in $1+1$ dimensions: an exponential growth, due to MI, from the unstable background $u_{0}(t)$, reaching the amplitude maximum in $(x_{0},t_{0})$, and the subsequent decay to the background, and they have been observed in many experiments; see, for instance: [14, 40, 70, 41, 50, 57]. See [26, 13] for the study of the linear instability of background solutions in coupled NLS equations.
[…]
@@
or, equivalently, the dependence on the extra space variables $y,z,\dots$ is slow. In this regime all the above MNLS equations are close to NLS, and one expects that suitable adiabatic deformations of the AB (2) and of the Peregrine instanton (3) be relevant. The goal of this paper is to uncover the universal analytic features of AWs of MNLS equations in the Q1D regime, concentrating our attentions on the first nonlinear stage of MI (NLSMI), and showing, at leading order, a universal behavior, independent of the particular model, and described in terms of suitable adiabatic deformations of the AB (2). Varying the initial data, the first NLSMI consists in various combinations of the following two basic processes. P1: the AW growth from the unstable background, followed by fission in the transversal slowly varying directions, with infinite speed at fission time. P2: the inverse process of fusion in the transversal slowly varying directions, with infinite speed at fusion time, followed by AW decay to the background. Fission and fusion are critical processes showing similarities to multidimensional wave breaking, and to phase transitions of second kind and critical exponent $1/2$. They may or may not coexist in the same nonlinear stage of MI, but while the universal process of “AW growth + fission” always occurs, fusion is less generic. We also show that, in the natural long wave limit, the above Q1D Akhmediev breather reduces to the Q1D analogue of the Peregrine AW, rationally localized in the space variables, also undergoing the above critical processes. At last we show, through numerical experiments on the HNLS equation, that the process of “AW growth + fission” is by no means restricted to the Q1D regime, extending to a finite region of the MI domain of the equation (see [22] for more details). The recurrence of Q1D AWs, another important feature of $x$-periodicity, turns out to be different for different MNLS models, and the quantitative description of these differences is postponed to the subsequent paper [24]. Due to the universality of the above processes, it is very plausible that they be observable in natural phenomena in which AWs of MNLS equations be relevant, like in water waves, nonlinear optics, plasma physics, and Bose-Einstein condensates.
@@ 2 AW fission and fusion
$\begin{array}[]{l}u_{lin}(x,Y,t)=e^{2it}\Big{[}1+{\epsilon}\Big{(}\frac{2|\alpha(Y)|}{\sigma(k_{x})}e^{\sigma(k_{x})t+i\phi}\cos[k_{x}(x-x_{1}(Y))]\\ +\frac{2|\beta(Y)|}{\sigma(k_{x})}e^{-\sigma(k_{x})t-i\phi}\cos[k_{x}(x-x_{0}(Y))]\Big{)}+O({\epsilon})\mbox{ oscillations}\Big{)}\Big{]},\end{array}$ (8)
@@
$\begin{array}[]{l}x_{1}(Y)=\frac{\arg(\alpha(Y))+\pi/2}{k_{x}},\ \ x_{0}(Y)=\frac{-\arg(\beta(Y))+\pi/2}{k_{x}},\\ \alpha(Y)=e^{-i\phi_{1}}c^{*}_{+}(Y)-e^{i\phi_{1}}c_{-}(Y),\ \ \beta(Y)=e^{i\phi_{1}}c^{*}_{-}(Y)-e^{-i\phi_{1}}c_{+}(Y).\end{array}$ (9)
@@
matching successfully with the following adiabatic deformation of the AB (2) [30]
$\begin{array}[]{l}u_{1}(x,Y,t)={\cal A}\left(x-x_{1}(Y),t-t_{1}(Y),\phi\right)e^{2it+2i\phi},\\ t_{1}(Y)=\frac{1}{\sigma(k_{x})}\log\left(\frac{\sigma^{2}}{2{\epsilon}|\alpha(Y)|}\right),\ \ |t-t_{1}(Y)|=O(1).\end{array}$ (11)
Equation (11) describes at leading order the first NLSMI, the same for equations (4) and for all MNLS equations, up to small corrections whose analytic investigation is postponed to a subsequent paper; therefore it describes the universal features of Q1D AWs. Also for the ENLS equation, well known to lead to a blow up of generic initial perturbations of $u_{0}$ [65], the first NLSMI is described by (11) if the MI time scale $\ln(1/{\epsilon})$ is much smaller than the Q1D scale $1/\delta$, implying that blow up occurs only after a number of recurrences.
To understand the main properties of the first NLSMI (11), we restrict the family of initial data in (7) to
@@
Since $\dot{x}_{1}(0)=\dot{t}_{1}(0)=0$, and $\ddot{t}_{1}(0)>0$, it follows that, at the max point $(x,Y)=(x_{10},0)$, ${u_{1}}_{Y}={u_{1}}_{x}=0$, and ${u_{1}}_{YY}(x_{1}(0),0,t)=-\ddot{t}_{1}(0){\cal A}_{t}(0,t-t_{0})$, $|t-t_{0}|\ll 1$. Consequently $(|u_{1}(x_{10},0,t)|^{2})_{YY}=-\ddot{t}_{1}(0)\rho_{t}(t)$, implying that $\mbox{sign}\!\left((|u_{1}(x_{10},0,t)|^{2})_{YY}\right)=\mbox{sign}(t-t_{0})$, $|t-t_{0}|\ll 1$. It follows that, when the Q1D AW reaches it max, its $Y$-curvature changes sign, from negative to positive sign, corresponding to fission in the $Y$ direction. Fission and fusion as critical processes. Fission and fusion are critical processes showing similarities i) to multidimensional wave breaking according to the integrable dispersionless Kadomtsev-Petviashvili (dKP) equation $(u_{t}+uu_{x})_{x}+u_{yy}=0$ (describing for instance the evolution of small amplitude, nearly one-dimensional waves in shallow water near the shore), and ii) to phase transitions of second kind and critical exponent $1/2$. Indeed, i) in the dKP wave breaking, a smooth localized initial datum undergoes the first wave breaking in a space point $(x_{b},y_{b})$ at time $t_{b}$, constructed from the initial data via the spectral transform developed for integrable dispersionless PDEs [46, 47]. For $0<t-t_{b}\ll 1$ the multivaluedness region develops from the breaking point $(x_{b},y_{b})$ with the law $\Delta x=O\left((t-t_{b})^{3/2}\right)$ and velocity $\Delta\dot{x}=O\left((t-t_{b})^{1/2}\right)$ (zero speed at breaking time) in the $x$ direction, like in the $1+1$ dimensional reduction $u_{t}+uu_{x}=0$ (the Hopf equation), while it develops with the law $\Delta y=O\left((t-t_{b})^{1/2}\right)$ and velocity $\Delta\dot{y}=O\left((t-t_{b})^{-1/2}\right)$ (infinite speed at breaking time) in the slowly varying direction [47, 48] (see Figure 4). ii) In the Landau-Ginzburg theory of phase transitions, the behavior of the order parameter around the critical temperature $T_{c}$ is a two-valued function proportional to $\pm\theta(T_{c}-T)\sqrt{T_{c}-T}$, where $\theta(\cdot)$ is the step function, and the susceptivity is singular at $T_{c}$, with critical exponent $1/2$ [54].
@@
Figure 7: The central figure shows the MI domain $0<k^{2}_{x}-k^{2}_{y}<4$ of the HNLS equation. The right figure shows four snapshots of the HNLS numerical evolution of a doubly periodic initial condition $1+{\epsilon}(e^{i\phi}\cos(k_{x}x)\cos(k_{y}y))$, with periods $L_{x}=4,\ L_{y}=10$, where $k_{x}=2\pi/L_{x},\ k_{y}=2\pi/L_{y}$ are the coordinates of the red points of the central figure, and $\phi=\arccos(\sqrt{k^{2}_{x}-k^{2}_{y}}/2)$; this AW undergoes fission. The left figure shows four snapshots of the HNLS numerical evolution for the same initial condition, but with different periods $L_{x}=2.2,\ L_{y}=2.89$, corresponding to the green points of the central figure; this AW does not undergo fission.
@@
Acknowledgments. F.C. and P.M.S. have been supported by the Research Project of National Interest PRIN2020 No. 2020X4T57A; F.C. was also supported by PRIN2022 No. 20223T577Z. We also acknowledge the useful support of the INFN Research Project MMNLP.